Solve.
If the hypotenuse of a right triangle is 12 feet and a side is 6 feet, what is the length of the other side. Express your answer in radical form.
a= 6 square root of 5
b= 3 square root of 6
c= 5 square root of 6
d= 6 square root of 3

Answer by checkley71(8403)

12^2=6^2+X^2
144=36+X^2
X^2=144-36
X^2=108
X=SQRT(36*3)
X=6SQRT3 ANSWER --D).